OpenStudy (anonymous):

A construction job will be finished in 4 days if 6 people are working. If 2 more people are added to the job, how many days will the job take to complete? MULTIPLE CHOICE ANSWERS: A)3 B)3.5 C)4 D)6

OpenStudy (anonymous):

Why won't anyone help? </3

hero (hero):

Looks like inverse proportional to me

OpenStudy (anonymous):

It's actually using measurements, but yes, it's used as a proportion. But i'm stuck. I can't get the proportion right. Can you help?

hero (hero):

I'll try

OpenStudy (anonymous):

Thank you :3

hero (hero):

Let's do it this way. Remember, in an "if, then" statement, if comes before then. In this case, "if x, then y", where x = people and y = days.

hero (hero):

so the beginning ratio is x:y = 6:4

hero (hero):

Applying that to the formula for inverse proportional, xy = k

hero (hero):

6*4 = k 24 = k

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

I figured the answer was A) But i'm probably wrong. </3

hero (hero):

so the only thing that changes is that we will add more people to the job. So instead of 6 people, there are 8. Now keeping k the same, we modify x to equal 8: 8y = 24 y = 24/8 y = 3

OpenStudy (anonymous):

That's what I thought too. Thanks! I think that should be the right answer.

OpenStudy (anonymous):

I have another one the same one as kthis, but i'm just so confused...

OpenStudy (anonymous):

But thanks! :D

hero (hero):

yw

hero (hero):

Are you going to post the next question? If so, close this one and then post the next one separately.